Given that this flight was neoy heavily delayed, find the probability that it travelled between 500 km and 5000 km
1
Expert's answer
2020-12-04T04:59:14-0500

Find condition probability:


P(between 500 km and 5000 km | not heavily delayed) =


= P(between 500 km and 5000 km AND not heavily delayed) / P( not heavily delayed) =


= [ P(between 500 km and 5000 km AND On time) + P(between 500 km and 5000 km AND Slightly delayed) ] / [P(On time) + P(Slightly delayed)]= (17+18) / (52 + 45) = 35/97 = 0.36